A 0.300 M solution of NaCl that contains 0.150 mole of NaCl would have a volume, in mL, of

Group of answer choices

5.00 × 103 mL

0.045 mL

5.00 × 102 mL

45.0 mL
Answer #1

Molecular weight of NaCl =(23+25.5g = 58.5g I mole Nace = 58.5 g 0.150 mole Nace =(58.5*0.150g = 8.775 g Now 1000 ml I (M) Na